OpenStudy (anonymous):

equation x^2+2x+3=0 roots are x1 and x2. the quadratic equation with the roots 1/x1 and 1/x2 is A. 3x^2+2x-1=0 B. 3x^2+2x+1=0 C. 3x^2 -2x+1=0 D. x^2-2x+3=0 E. x^2-2x-3=0

OpenStudy (vishweshshrimali5):

Is you remember the formula I used in previous question: If u and v are the roots of a quadratic equation ax^2 + bx + c = 0, then, \[\large{u+v = \cfrac{-b}{a}}\] \[\large{uv = \cfrac{c}{a}}\]

OpenStudy (vishweshshrimali5):

Now, here the roots of the equation x^2 + 2x + 3 = 0 are x1 and x2. So, can you calculate x1 + x2 and x1 * x2 ?

OpenStudy (vishweshshrimali5):

@dinisha you there ?

OpenStudy (anonymous):

ya i am trying to solve

OpenStudy (vishweshshrimali5):

Ok take your time :)

OpenStudy (anonymous):

u+v=-2/1=-2 uv=3/1=3 x1+x2=-2 x1*x2=3

OpenStudy (vishweshshrimali5):

Very good

OpenStudy (vishweshshrimali5):

Now, you have been asked to find an equation with roots 1/x1 and 1/x2. Okay let the equation be : ax^2 + bx + c = 0 Now, you know that this equation has roots 1/x1 and 1/x2. (1) Find out 1/x1 + 1/x2 and 1/x1 * 1/x2 using the values of x1 + x2 and x1*x2. (2) Find out 1/x1 + 1/x2 and 1/x1 * 1/x2 in terms of a, b and c.

OpenStudy (anonymous):

in that case if 1/x1+1/x2 will it be -1/2 and 1/x1*1/x2=1/3 ???

OpenStudy (vishweshshrimali5):

1/x1 * 1/x2 is correct. But, (1/x1) + (1/x2) is not. See: \[\large{\cfrac{1}{x_1} + \cfrac{1}{x_2} = \cfrac{x_1+x_2}{x_1\times x_2}}\]

OpenStudy (vishweshshrimali5):

Now, you already know the values of x1 + x2 and x1*x2.

OpenStudy (vishweshshrimali5):

@dinisha ?

OpenStudy (anonymous):

is it -2/3 ??? @vishweshshrimali5

OpenStudy (vishweshshrimali5):

Yes

OpenStudy (vishweshshrimali5):

Now tell me what is 1/x1 + 1/x2 and 1/x1 * 1/x2 in terms of a,b and c.

OpenStudy (anonymous):

wait a minute i am trying now!!

OpenStudy (vishweshshrimali5):

take your time :)

OpenStudy (anonymous):

1/x1+1/x2 = -a/b 1/x1*1/x2 = -b/c

OpenStudy (vishweshshrimali5):

Nope

OpenStudy (vishweshshrimali5):

Some of roots (for ax^2 + bx + c = 0) = -b/a Product of roots = c/a

OpenStudy (anonymous):

a=1, b=-2, c=3 ax^2+bx+c=0 x^2-2x+3=0 Am I right?

OpenStudy (vishweshshrimali5):

Hey hey slow down... First of all how did you obtain a,b, and c ? Second thing that's wrong.

OpenStudy (anonymous):

whr did i went wrong??

OpenStudy (vishweshshrimali5):

First of all tell me how did you find out a, b, and c ?

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 Some of roots (for ax^2 + bx + c = 0) = -b/a Product of roots = c/a \(\color{blue}{\text{End of Quote}}\) Thus, 1/x1 + 1/x2 = -b/a and 1/x1 * 1/x2 = c/a

OpenStudy (anonymous):

yaa soo wat i did was i compared the equation with terms and the answer i got for 1/x1+1/x2 and 1/x2*1/x2

OpenStudy (vishweshshrimali5):

Okay lets see: 1/x1 + 1/x2 = -2/3 = -b/a 1/x1 * 1/x2 = 1/3 = c/a So, a = 3, b= 2, c = 1 So, ax^2 + bx + c = 0 => 3x^2 + 2x + 1 = 0

OpenStudy (vishweshshrimali5):

Did you get your mistake ?

OpenStudy (vishweshshrimali5):

@dinisha ??

OpenStudy (anonymous):

yaaa i got it!! thank u :) @vishweshshrimali5

OpenStudy (vishweshshrimali5):

np :)